# Tacit Learning with Adaptive Information Selection for Cooperative Multi-Agent Reinforcement Learning

### v1

@@ Abstract.
Multi-agent Reinforcement Learning, Adaptive Information Selection, Tacit Learning
@@ 2. Related Work
Continous Communication CTDE methods typically incorporate global information solely in the mixing network, with agents limited to observing only local information. To augment the decision-making capabilities of agents, numerous studies have adopted communication methods to introduce global information to the agents. One pioneering work in this domain is RIAL with DIAL Foerster et al. (2016), which aims to facilitate communication in CTDE through predefined topological structures. Meanwhile, CommNet was introduced as the first differentiable framework in MARL Sukhbaatar et al. (2016). Subsequent works, such as ATOC Jiang and Lu (2018), I2C Ding et al. (2020)), and DHCG Sukhbaatar et al. (2016), utilize state-dependent communication graphs to address these challenges. ATOC and I2C also introduce gating mechanisms to regulate communication links between agents. Recently, novel communication methods such as TarMAC Das et al. (2018)and NDQ Wang et al. (2019a)have been proposed. TarMAC employs multi-layer attention modules for multi-round communication to strengthen connections between agents. NDQ aims to reduce communication overhead by learning nearly decomposable value functions.
@@ 6. Limitations
In this section, we examine three potential limitations of SICA. First, the Regeneration Block may encounter difficulties when there is no correlation between the trajectories or observations of the agents, as it relies on similar historical information for regeneration. Although we mitigate this limitation by incorporating observations from other agents into a mini buffer, there may still be instances where those agents fail to utilize this portion of historical data. Second, while we are confident in SICA’s performance in large-scale multi-agent environments, our experiments are constrained by computational resources, necessitating further evaluation of its effectiveness in larger-scale scenarios. Thirdly, when applying SICA to new tasks, adjustments to multiple hyperparameters will be necessary. Finally, we implemented parameter sharing among the agents to accelerate the training process, as not using this technique would significantly increase the training time.

### v2

@@ Abstract.
Multi-agent Reinforcement Learning, Tacit Learning, Adaptive Information Selection
@@ 2. Related Work
Continous Communication CTDE methods typically incorporate global information solely in the mixing network, with agents limited to observing only local information. To augment the decision-making capabilities of agents, numerous studies have adopted communication methods to introduce global information to the agents. One pioneering work in this domain is RIAL with DIAL Foerster et al. (2016), which aims to facilitate communication in CTDE through predefined topological structures. Meanwhile, CommNet was introduced as the first differentiable framework in MARL Sukhbaatar et al. (2016). Subsequent works, such as ATOC Jiang and Lu (2018) and I2C Ding et al. (2020), utilize state-dependent communication graphs to address these challenges. ATOC and I2C also introduce gating mechanisms to regulate communication links between agents. Recently, novel communication methods such as TarMAC Das et al. (2018) and NDQ Wang et al. (2019a) have been proposed. TarMAC employs multi-layer attention modules for multi-round communication to strengthen connections between agents. NDQ aims to reduce communication overhead by learning nearly decomposable value functions.
@@ 6. Limitations
In this section, we discuss four limitations of SICA. First, the Regeneration Block may struggle when agents’ observations or trajectories are dissimilar, although a mini-buffer helps alleviate this issue. Second, our experiments were constrained by computational resources, and further testing in larger environments is needed. Third, applying SICA to new tasks requires tuning several hyperparameters. Finally, parameter sharing is used to accelerate training; without it, training times would be significantly longer.